xXAikoXx:

Math question (8th grade)

xXAikoXx:

ihatelifeFS420:

u do 28*8=224 then u find out what times 20 equals 224

ihatelifeFS420:

@astrid1 am I right

surjithayer:

\[\frac{ t }{ 28}=\frac{ 20 }{ 8 }\] \[t=\frac{ 20 }{ 8 }\times 28=70\]

tahyla:

70

AnimeWeeebLMAO:

This is easy, T/28 obviously equals 20/8. So that means also equals 20/8, multiply that by 28 and you get 70.

xXAikoXx:

@animeweeeblmao wrote:
This is easy, T/28 obviously equals 20/8. So that means also equals 20/8, multiply that by 28 and you get 70.
Thanks